# The Atacama Cosmology Telescope: DR6 Gravitational Lensing and SDSS BOSS cross-correlation measurement and constraints on gravity with the EG statistic

### v1

@@ VII.3 Sensitivity to analysis choices
To evaluate the robustness of the presented results (Section VII.2) we investigate the sensitivity of the results to several analysis choices. These analysis choices were fixed before unblinding the data and the sensitivity to these analysis choices was only evaluated after the baseline results were finalized. See Section V for a detailed description of the blinding approach.
@@
In this analysis, we report combined constraints from the SDSS BOSS North and South patches after testing the angular cross-power spectra for each region for consistency (see Section VI.2.6). For completeness, we also show the results for the angular power spectra calculated only for the North or South patches. Note that the values for the full maps are used, as the BOSS data vectors for the RSD analysis are only available for the full map [26]. Except for , the two measurements are independent and therefore can differ statistically. We find them consistent with expectations.
For the baseline analysis, the full galaxy maps are used to calculate . In Section VII.1.2 we evaluated the angular auto-power spectrum for consistency between the full maps and the overlap to estimate a systematic error budget. We here for completeness show the result for the alternative analysis choice of restricting the to the footprint of the ACT DR6 maps. As the two measurements are consistent the resulting values are also close consistent.
Apart from these analysis choices, we can also evaluate the impact of parts of the analysis. We correct for the effect of magnification bias in our measurement (Section VI.1.5). Below the gray dashed line in Fig. 17 the results for without this correction are shown. For ACT + LOWZ the magnification bias effect is negligible. For ACT + CMASS the correction is 2–3%, a small difference compared to the statistical uncertainty. The overall conclusions of the analysis are not affected by the magnification bias correction. For future surveys with larger statistical constraining power, the correction will become increasingly impactful.
@@ VII.4 Combining ACT and Planck
In Section VI.4 we discussed in detail how we calculated the minimum variance combination of the angular cross-power spectra at the bandpower level, , and estimated the covariance for the combination. To estimate this combination numerically accurate we here switch to using the analytic covariance matrices that we have confirmed to give consistent results (see Sections VI.1.3 and VII.3).
@@ VII.5 Results in context
The ACT DR6 lensing map used in this analysis gives constraints on independent of Planck. In Fig. 20 the Planck based constraints can be compared to the ACT + CMASS [Orange circle] and ACT + LOWZ [Orange pentagon] constraints. Both measurements match the consistently analyzed Planck results from Wenzl et al. [26] within the confidence range. Furthermore, the cosmological interpretation of both measurements is also consistent with neither showing statistical evidence for tension with the CDM prediction. The analysis is overall limited by the uncertainty of the measurement, making this a strong consistency check between the Planck and ACT CMB lensing-based results. We can further compare the results to previous analyses using Planck CMB lensing data. The result from Singh et al. [32] using Planck + LOWZ is consistent within the confidence range and the Planck + CMASS results differ by slightly more than the confidence range. We note that if we assumed Gaussianity for the result as done in Singh et al. [32] then both measurements would match within the confidence range (see the last line of Fig. 17). The ACT + CMASS measurement result is approximately in between the low Planck + CMASS measurement reported in Pullen et al. [31] and the other published measurements [32, 26]. It does not disagree significantly with any of the three measurements. Notably, however, the result is not in tension with the CDM predictions similar to [32, 26] whereas the Pullen et al. [31] result does show a significant tension with the prediction.
@@ VIII Conclusion
Based on the new constraints on the cross-correlation we derived estimates of the statistic of gravity. To estimate the statistic we additionally used previous analyses of the SDSS BOSS samples, specifically we used the galaxy auto-correlation and results reported for CMASS and LOWZ in Wenzl et al. [26]. The resulting values for were summarized in Table 2. We tested our measurement both for consistency with scale independence as predicted by GR generally and tested the overall constraint on against the predicted values assuming CDM based on fits to primary CMB anisotropies and BAO. We found both the ACT only and combined constraints statistically consistent with the predictions for our fiducial cosmology. This indicates that within our measurement uncertainty, the observed growth of structure is consistent with predictions using CDM fitted to the background expansion and early universe. Our combined measurements achieve fractional errors of approximately 15% with ACT and Planck CMB lensing together with CMASS and 23% with LOWZ. Our results are in line with a wide range of investigations of in the literature most of which found similar consistency with GR predictions [for a recent overview see 33]. The developed techniques to measure are designed to be unbiased to a much higher level with systematic errors of only 3% for ACT + CMASS and 4% for ACT + LOWZ. This systematic error is dominated by the conservative accounting for a potential effective bias difference between the full galaxy maps and the overlap with ACT which could be avoided in future work by recalculating the RSD analysis for the specific overlaps with CMB lensing maps. The main astrophysical systematic of galaxy bias evolution with redshift can also be mitigated further for analyses with larger datasets by calculating in narrower bins in redshift. Therefore future data with larger statistical constraining power offers the opportunity to improve upon these constraints and probe gravity at the percent level. This will allow discriminating even smaller deviations from the predictions for the growth of the structure in alternative gravitational scenarios compared to a cosmological constant [28].
@@ Acknowledgements.
We thank Martin White for helpful discussions on best practices for HODs from SDSS BOSS publications. We thank Sigurd K. Næss and Bruce Partridge for helpful comments on a draft of this paper. The work of LW and RB is supported by NSF grant AST-2206088 and NASA ROSES grant 12-EUCLID12-0004. SC acknowledges the support of the National Science Foundation at the Institute for Advanced Study. GSF acknowledges support through the Isaac Newton Studentship and the Helen Stone Scholarship at the University of Cambridge. GSF, FJQ, BS, and NM acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 851274). GAM is part of Fermi Research Alliance, LLC under Contract No. DE-AC02-07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. MM acknowledges support from NSF grants AST-2307727 and AST-2153201 and NASA grant 21-ATP21-0145. AvE acknowledges support from NASA grants 80NSSC23K0747 and 80NSSC23K0464. CS acknowledges support from the Agencia Nacional de Investigación y Desarrollo (ANID) through Basal project FB210003. IH acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 849169). OD acknowledges support from a SNSF Eccellenza Professorial Fellowship (No. 186879). EC acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 849169). Support for ACT was through the U.S. National Science Foundation through awards AST-0408698, AST-0965625, and AST-1440226 for the ACT project, as well as awards PHY-0355328, PHY-0855887 and PHY-1214379. Funding was also provided by Princeton University, the University of Pennsylvania, and a Canada Foundation for Innovation (CFI) award to UBC. ACT operated in the Parque Astronómico Atacama in northern Chile under the auspices of the Agencia Nacional de Investigación y Desarrollo (ANID). The development of multichroic detectors and lenses was supported by NASA grants NNX13AE56G and NNX14AB58G. Detector research at NIST was supported by the NIST Innovations in Measurement Science program. Computing for ACT was performed using the Princeton Research Computing resources at Princeton University, the National Energy Research Scientific Computing Center (NERSC), and the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by the CFI under the auspices of Compute Canada, the Government of Ontario, the Ontario Research Fund–Research Excellence, and the University of Toronto. We thank the Republic of Chile for hosting ACT in the northern Atacama, and the local indigenous Licanantay communities whom we follow in observing and learning from the night sky. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC award HEP-ERCAPmp107 The sky simulations used in this paper were developed by the WebSky Extragalactic CMB Mocks team, with the continuous support of the Canadian Institute for Theoretical Astrophysics (CITA), the Canadian Institute for Advanced Research (CIFAR), and the Natural Sciences and Engineering Council of Canada (NSERC), and were generated on the Niagara supercomputer at the SciNet HPC Consortium (cite https://arxiv.org/abs/1907.13600). SciNet is funded by: the Canada Foundation for Innovation under the auspices of Compute Canada; the Government of Ontario; Ontario Research Fund - Research Excellence; and the University of Toronto. SDSS-III is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS-III Collaboration including the University of Arizona, the Brazilian Participation Group, Brookhaven National Laboratory, Carnegie Mellon University, University of Florida, the French Participation Group, the German Participation Group, Harvard University, the Instituto de Astrofisica de Canarias, the Michigan State/Notre Dame/JINA Participation Group, Johns Hopkins University, Lawrence Berkeley National Laboratory, Max Planck Institute for Astrophysics, Max Planck Institute for Extraterrestrial Physics, New Mexico State University, New York University, Ohio State University, Pennsylvania State University, University of Portsmouth, Princeton University, the Spanish Participation Group, University of Tokyo, University of Utah, Vanderbilt University, University of Virginia, University of Washington, and Yale University.Appendix A Combining the cross-correlation measurements
@@
In this section, we describe how we calculate the combined angular cross-power spectrum of the measurements based on ACT SDSS and Planck SDSS. We combine the measurements at the bandpower level using inverse variance weighting. Importantly the ACT and Planck CMB lensing maps overlap and therefore, the cross-correlation measurements have a non-negligible correlation. In the overlap region between the Planck and ACT CMB lensing maps the cosmic variance is correlated and both cross-correlation measurements use the same galaxy samples. We need to account for this correlation when estimating the covariance of the combined measurement.

### v2

@@ VI.3 Sensitivity to analysis choices
To evaluate the robustness of the presented results (Section VI.2) we investigate the sensitivity of the results to several analysis choices. These analysis choices were fixed before unblinding the data and the sensitivity to these analysis choices was only evaluated after the baseline results were finalized. See Section IV for a detailed description of the blinding approach.
@@
In this analysis, we report combined constraints from the SDSS BOSS North and South patches after testing the angular cross-power spectra for each region for consistency (see Section V.2.6). For completeness, we also show the results for the angular power spectra calculated only for the North or South patches. Note that the values for the full maps are used, as the BOSS data vectors for the RSD analysis are only available for the full map [27]. Except for , the two measurements are independent and therefore can differ statistically. We find them consistent with expectations.
For the baseline analysis, the full galaxy maps are used to calculate . In Section VI.1.2 we evaluated the angular auto-power spectrum for consistency between the full maps and the overlap to estimate a systematic error budget. We here for completeness show the result for the alternative analysis choice of restricting the to the footprint of the ACT DR6 maps. As the two measurements are consistent the resulting values are also closely consistent. For additional discussion of the measurements used as part of the estimate, we refer the interested reader to Wenzl et al. [27].
Apart from these analysis choices, we can also evaluate the impact of parts of the analysis. We correct for the effect of magnification bias in our measurement (Section V.1.4). Below the gray dashed line in Fig. 14 the results for without this correction are shown. For ACT + LOWZ the magnification bias effect is negligible. For ACT + CMASS the correction is 2–3%, a small difference compared to the statistical uncertainty. The overall conclusions of the analysis are not affected by the magnification bias correction. For future surveys with larger statistical constraining power, the correction will become increasingly impactful.
@@ VI.4 Combining ACT and Planck
In Section V.4 we discussed in detail how we calculated the minimum variance combination of the angular cross-power spectra at the bandpower level, , and estimated the covariance for the combination. To estimate this combination numerically accurate we here switch to using the analytic covariance matrices that we have confirmed to give consistent results (see Sections V.1.2 and VI.3).
@@ VI.5 Results in context
The ACT DR6 lensing map used in this analysis gives constraints on independent of Planck. In Fig. 17 the Planck based constraints can be compared to the ACT + CMASS [Orange circle] and ACT + LOWZ [Orange pentagon] constraints. Both measurements match the consistently analyzed Planck results from Wenzl et al. [27] within the confidence range. Furthermore, the cosmological interpretation of both measurements is also consistent with neither showing statistical evidence for tension with the CDM prediction. The analysis is overall limited by the uncertainty of the measurement, making this a strong consistency check between the Planck and ACT CMB lensing-based results. We can further compare the results to previous analyses using Planck CMB lensing data. The result from Singh et al. [33] using Planck + LOWZ is consistent within the confidence range and the Planck + CMASS results differ by slightly more than the confidence range. We note that if we assumed Gaussianity for the result as done in Singh et al. [33] then both measurements would match within the confidence range (see the last line of Fig. 14). The ACT + CMASS measurement result is approximately in between the low Planck + CMASS measurement reported in Pullen et al. [32] and the other published measurements [33, 27]. It does not disagree significantly with any of the three measurements. Notably, however, the result is not in tension with the CDM predictions similar to [33, 27] whereas the Pullen et al. [32] result does show a larger difference with the prediction.
@@ VII Conclusion
Based on the new constraints on the cross-correlation we derived estimates of the statistic of gravity. To estimate the statistic we additionally used previous analyses of the SDSS BOSS samples, specifically we used the galaxy auto-correlation and results reported for CMASS and LOWZ in Wenzl et al. [27]. The resulting values for were summarized in Table 2. We tested our measurement both for consistency with scale independence as predicted by GR generally and tested the overall constraint on against the predicted values assuming CDM based on fits to primary CMB anisotropies and BAO. We found both the ACT only and combined constraints statistically consistent with the predictions for our fiducial cosmology. This indicates that within our measurement uncertainty, the observed growth of structure is consistent with predictions using CDM fitted to the background expansion and early universe. Our combined measurements achieve fractional errors of approximately 15% with ACT and Planck CMB lensing together with CMASS and 23% with LOWZ. Our results are in line with a wide range of investigations of in the literature most of which found similar consistency with GR predictions [for a recent overview see 34]. The developed techniques to measure are designed to be unbiased to a much higher level with systematic errors of only 3% for ACT + CMASS and 4% for ACT + LOWZ. This systematic error is dominated by the conservative accounting for a potential effective bias difference between the full galaxy maps and the overlap with ACT which could be avoided in future work by recalculating the RSD analysis for the specific overlaps with CMB lensing maps. The main astrophysical systematic of galaxy bias evolution with redshift can also be mitigated further for analyses with larger datasets by calculating in narrower bins in redshift. Therefore future data with larger statistical constraining power offers the opportunity to improve upon these constraints and probe gravity at the percent level. This will allow discriminating even smaller deviations from the predictions for the growth of the structure in alternative gravitational scenarios and General Relativity [29].
@@ Acknowledgements.
We thank Martin White for helpful discussions on best practices for HODs from SDSS BOSS publications. We thank Sigurd K. Næss and Bruce Partridge for helpful comments on a draft of this paper. The work of LW and RB is supported by NSF grant AST-2206088 and NASA ROSES grant 12-EUCLID12-0004. SC acknowledges the support of the National Science Foundation at the Institute for Advanced Study through NSF/PHY 2207583. GSF acknowledges support through the Isaac Newton Studentship and the Helen Stone Scholarship at the University of Cambridge. GSF, FJQ, BS, and NM acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 851274). GAM is part of Fermi Research Alliance, LLC under Contract No. DE-AC02-07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. MM acknowledges support from NSF grants AST-2307727 and AST-2153201 and NASA grant 21-ATP21-0145. AvE acknowledges support from NASA grants 80NSSC23K0747 and 80NSSC23K0464. CS acknowledges support from the Agencia Nacional de Investigación y Desarrollo (ANID) through Basal project FB210003. IH acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 849169). OD acknowledges support from a SNSF Eccellenza Professorial Fellowship (No. 186879). EC acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 849169). Support for ACT was through the U.S. National Science Foundation through awards AST-0408698, AST-0965625, and AST-1440226 for the ACT project, as well as awards PHY-0355328, PHY-0855887 and PHY-1214379. Funding was also provided by Princeton University, the University of Pennsylvania, and a Canada Foundation for Innovation (CFI) award to UBC. ACT operated in the Parque Astronómico Atacama in northern Chile under the auspices of the Agencia Nacional de Investigación y Desarrollo (ANID). The development of multichroic detectors and lenses was supported by NASA grants NNX13AE56G and NNX14AB58G. Detector research at NIST was supported by the NIST Innovations in Measurement Science program. Computing for ACT was performed using the Princeton Research Computing resources at Princeton University, the National Energy Research Scientific Computing Center (NERSC), and the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by the CFI under the auspices of Compute Canada, the Government of Ontario, the Ontario Research Fund–Research Excellence, and the University of Toronto. We thank the Republic of Chile for hosting ACT in the northern Atacama, and the local indigenous Licanantay communities whom we follow in observing and learning from the night sky. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC award HEP-ERCAPmp107 The sky simulations used in this paper were developed by the WebSky Extragalactic CMB Mocks team, with the continuous support of the Canadian Institute for Theoretical Astrophysics (CITA), the Canadian Institute for Advanced Research (CIFAR), and the Natural Sciences and Engineering Council of Canada (NSERC), and were generated on the Niagara supercomputer at the SciNet HPC Consortium (cite https://arxiv.org/abs/1907.13600). SciNet is funded by: the Canada Foundation for Innovation under the auspices of Compute Canada; the Government of Ontario; Ontario Research Fund - Research Excellence; and the University of Toronto. SDSS-III is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS-III Collaboration including the University of Arizona, the Brazilian Participation Group, Brookhaven National Laboratory, Carnegie Mellon University, University of Florida, the French Participation Group, the German Participation Group, Harvard University, the Instituto de Astrofisica de Canarias, the Michigan State/Notre Dame/JINA Participation Group, Johns Hopkins University, Lawrence Berkeley National Laboratory, Max Planck Institute for Astrophysics, Max Planck Institute for Extraterrestrial Physics, New Mexico State University, New York University, Ohio State University, Pennsylvania State University, University of Portsmouth, Princeton University, the Spanish Participation Group, University of Tokyo, University of Utah, Vanderbilt University, University of Virginia, University of Washington, and Yale University.Appendix A Combining the cross-correlation measurements
@@
Our procedure to calculate the combined angular cross-power spectrum of the measurements based on ACT SDSS and Planck SDSS was as follows. We combine the measurements at the bandpower level using inverse variance weighting. Importantly the ACT and Planck CMB lensing maps overlap and therefore, the cross-correlation measurements have a non-negligible correlation. In the overlap region between the Planck and ACT CMB lensing maps the cosmic variance is correlated and both cross-correlation measurements use the same galaxy samples. We need to account for this correlation when estimating the covariance of the combined measurement.